What is the next step in the problem 6 1/3 - 2 2/5 ?

Answer by friesr(113)

you need to get a common denominator. I also like to make them top heavy fractions and reduce them later.
-
6 1/3 = 19/3 and 2 2/5 = 12/5
-
if we multiply 19/3 by 5/5 we get 95/15
-
if we multiply 12/5 by 3/3 we get 36/15
-
95/15 - 36/15 = 59/15
-
Reduced down we get 3 14/15